# On kinks and solitons in discrete nonlinear transmission line

### 17 Jan 2024 v2

@@ Abstract
We studied lossless discrete transmission lines constructed from ideal linear inductors and nonlinear capacitors. The travelling waves with finite support in such systems are the kinks and the solitons. We calculated the speeds and the profiles both of the kinks and of the solitons. We also studied dissipative kinks in the lossy transmission lines, with ohmic resistors shunting the inductors. The losses influence the dissipative kink speed and profile.
@@ II.1 The travelling waves
We take the voltages $V_{n}$ and the currents which pass though the inductors $I_{n}$ as the dynamical variables. The circuit equations are
@@
$L\frac{dI^{c}_{n}}{dt}$ $=V_{n+1}-2V_{n}+V_{n-1},$ (1b)
where $Q_{n}=Q(V_{n})$ and the nonlinear nature of the transmission line is determined by the function $Q(V)$. We also took into account the connection between the current through the capacitor $dI^{c}_{n}$ and the currents between the inductors
$I^{c}_{n}=I_{n-1,n}-I_{n,n+1}.$ (2)
Let us write down Eq. (II.1) in the quasi-continuum approximation, which was proposed in our previous publications kogan2 ; kogan3 . The approximation consists in treating $n$ as the continuous variable $Z$, expanding the r.h.s. of the equation (1b) in Taylor series with respect to $Z$ and keeping only the first two terms of the expansion
$V_{n+1}-2V_{n}+V_{n-1}=\frac{\partial^{2}V}{\partial Z^{2}}+\frac{1}{12}\frac{\partial^{4}V}{\partial Z^{4}}.$ (3)
After that (II.1) takes the form
$\frac{\partial Q}{\partial t}$ $=I^{c},$ (4a)
$L\frac{\partial I^{c}}{\partial t}$ $=-\frac{\partial}{\partial Z}\left(\frac{\partial V}{\partial Z}+\frac{1}{12}\frac{\partial^{3}V}{\partial Z^{3}}\right).$ (4b)
In this paper we are interested in the travelling wave solutions of (II.1), that is the dependence of all the variables upon $Z$ and $t$ is of the form
@@
where $U$ is the speed of the travelling wave (the distance is measured in the units of the transmission line spacial period $\Lambda$). This gives us the opportunity to perform integration in (4b) to get
$ULI^{c}=\frac{\partial V}{\partial Z}+\frac{1}{12}\frac{\partial^{3}V}{\partial Z^{3}}.$ (6)
@@
$LU^{2}\frac{dQ}{dZ}=\frac{dV}{dZ}+\frac{1}{12}\frac{d^{3}V}{dZ^{3}}.$ (7)
An obvious way to integrate (7) starts by going to the second order ODE
$LU^{2}Q=V+\frac{1}{12}\frac{d^{2}V}{dZ^{2}}+\text{const}.$ (8)
However in this paper we’ll go along the different way. Presenting $dQ/dz$ as $dQ/dz=(dQ/dV)(dV/dZ)$, introducing new dependent variable
@@
$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}=LU^{2}\frac{dQ(V)}{dV}-1.$ (11)
@@
$LU^{2}\left(\frac{dQ}{dV}\right)_{V=V_{1,2}}-1>0.$ (13)
The inequality (13) has a simple physical meaning. Consider the ”sound” waves that is small amplitude, smooth waves propagating on the homogeneous backgrounds $V=V_{1,2}$. These waves are described by the equation obtained by linearizing (7) near $V=V_{1,2}$ and ignoring the term with the third order derivative
$Lu^{2}\left(\frac{dQ}{dV}\right)_{V=V_{1,2}}\frac{dV}{dZ}=\frac{dV}{dZ}.$ (14)
Hence the speed of the sound is
$\frac{1}{Lu_{1,2}^{2}}=\left(\frac{dQ}{dV}\right)_{V=V_{1,2}}.$ (15)
@@
From (10a) follows
$\left(\frac{dE^{2}}{dV}\right)_{V=V_{1,2}}=0.$ (18)
The conditions (17) and (18) mean that the particle is at rest at the origin at ”time” equal to $V_{1}$ and then again at ”time” equal to $V_{2}$.
@@
$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}=LU^{2}\left(\frac{dQ}{dV}\right)_{V=V_{1,2}}-1=\text{const}.$ (19)
Taking into account the boundary conditions (17) and (18) we obtain
$E^{2}\sim\left(V-V_{1,2}\right)^{2},\hskip 28.45274pt\text{ in the vicinity of }V_{1,2}.$ (20)
@@
$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}=LU^{2}(\alpha+2\beta V+3\gamma V^{2})-1.$ (22)
For brevity we’ll call Eq. (21) and (22) the cubic case.
@@ II.2 The kinks
$V(Z)=\frac{V_{2}+V_{1}}{2}+\frac{V_{2}-V_{1}}{2}\tanh\left[\frac{V_{2}-V_{1}}{2}\psi Z\right]$ (29)
(to satisfy the boundary conditions we have to choose positive value of $\psi$). After finding $V(Z)$ we can from (1b) obtain the current
$I(Z)=\frac{\psi\left(V_{2}-V_{1}\right)}{2LU}\tanh\left[\frac{V_{2}-V_{1}}{2}\psi Z\right].$ (30)
@@ II.3 The solitons
With three boundary conditions on $E(V)$ supplied by (17), (18) and (31) for the soliton problem of the three parameters $V_{1}V_{3},U^{2}$ only two (say $V_{1}$ and $V_{3}$) can be considered as free parameters. Multiplying both sides of (7) by $V_{3}-V$, integrating the equation with respect to $V$ from $V_{1}$ to $V_{3}$ and taking into account the boundary conditions we obtain the speed of the soliton
@@
From (II.3) we see that the soliton exists both for positive and negative $\beta$. For positive $\beta$ we have dark soliton ($V_{3}<V_{1}$), and for negative $\beta$ – bright soliton ($V_{3}>V_{1}$). From (38c) we see that for any sign of $\beta$ the soliton speed is larger than the sound speed at the background, given by equation
$\frac{1}{LC_{0}u^{2}}=\alpha+2\beta V_{1},$ (39)
in line with our previous general statement that the soliton is supersonic.
@@
Similar to what was done in the previous subsection, we can use (1b) to find the current and obtain
$I(Z)=\frac{\phi\left(V_{3}-V_{1}\right)}{2LU}\operatorname{sech}^{2}\left(\frac{V_{3}-V_{1}}{2}\phi Z\right).$ (42)
@@ II.4 The quasi-continuum approximation
We would like to discuss an opportunity (and motivation) to go beyond the limits of the quasi-continuum approximation used in the paper. To do it, one may keep the third term in the Taylor expansion of the discrete second order derivative in the r.h.s. of (1b) kogan2 ; kogan3
@@
$LU^{2}\frac{dQ}{dZ}=\frac{dV}{dZ}+\frac{1}{12}\frac{d^{3}V}{dZ^{3}}+\frac{1}{360}\frac{d^{5}V}{dZ^{5}}.$ (44)
@@ III The dissipative kinks
In this case (2) is modified to
@@
hence (II.1) takes the form
$\frac{\partial Q}{\partial t}$ $=I^{c},$ (46a)
$L\frac{\partial I^{c}}{\partial t}$ $=-\frac{\partial}{\partial Z}\left(\frac{\partial V}{\partial Z}+\frac{1}{12}\frac{\partial^{3}V}{\partial Z^{3}}+\frac{L}{R}\frac{\partial^{2}V}{\partial t\partial Z}\right).$ (46b)
@@
After repeating exactly what was done in Section II.1 we obtain in a modified form Eq. (11)
$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}+\frac{LU}{R}\frac{dE}{dV}=LU^{2}\frac{dQ(V)}{dV}-1$ (47)
and (in the cubic case) Eq. (22)
$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}+\frac{LU}{R}\frac{dE}{dV}=LU^{2}(\alpha+2\beta V+3\gamma V^{2})-1.$ (48)
We again impose on $V(Z)$ the boundary conditions (12), which lead to the boundary conditions (17) and (18) for $E(V)$. Equation (47) in the vicinity of $V_{1,2}$ has the form
$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}+\frac{LU}{R}\frac{dE}{dV}=\text{const},$ (49)
@@
In the cubic case we regain Eq. (25), and there exists kogan4 a simple polynomial solution of (48) in the form (26), which we repeat here
$E(V)=\psi(V_{1}-V)(V-V_{2}),$ (50)
where $\psi$ is a constant. Substituting (50) into (48) and equating the coefficients before the same powers of $V$ we obtain after simple algebra the generalization of (II.2)
@@
$V_{1}+V_{2}$ $=-\frac{2\beta}{3\gamma}+\frac{4LU}{\psi R}..$ (51b)
@@
Equation (50) shows that in the lossy line the kinks are of the same shape as in the lossless line (given by (29)). However, the value of $V_{2}$ (for a given $V_{1}$) depends upon $R$ (Eq. (51b) instead of (27b)) and so does the kink speed, in spite of the validity in both cases of Eq. (25).
There exists different from (50) type of solutions of (47) (hence different from (29) type of solutions of (III), similar to the shock waves we considered previously kogan1 ; kogan2 ; kogan3 ; kogan4 . In the present paper we don’t consider such solutions.

### 24 Jan 2024 v3

# On the kinks, the solitons and the shocks in discrete nonlinear transmission line

@@ Abstract
We studied lossless discrete transmission lines constructed from ideal linear inductors and nonlinear capacitors. The travelling waves with finite support in such systems are the kinks and the solitons, which speeds and profiles were calculated. We also studied lossy transmission lines. The travelling waves with finite support in such systems are the dissipative kinks and the shocks, which speeds and profiles were also calculated.
@@ II.1 The travelling waves
We take the voltages $V_{n}$ and the currents which pass though the inductors $I_{n}$ as the dynamical variables. The circuit equations (Kirchhoff laws) are
@@
$L\frac{dI_{n,n+1}}{dt}$ $=V_{n}-V_{n+1},$ (1b)
$I^{c}_{n}$ $=I_{n-1,n}-I_{n,n+1}.$ (1c)
where $Q_{n}=Q(V_{n})$ and the nonlinear nature of the transmission line is determined by the (monotonically increasing) function $Q=Q(V)$. We can exclude $I_{n,n+1}$ from (1b) and (1c) and present these two equations a single one
$L\frac{dI^{c}_{n}}{dt}=V_{n+1}-2V_{n}+V_{n-1}.$ (2)
Let us write down Eq. (2) in the quasi-continuum approximation, which was proposed in our previous publications kogan2 ; kogan3 . The approximation consists in treating $n$ as the continuous variable $Z$, expanding the r.h.s. of the equation (2) in Taylor series with respect to $Z$ and keeping only the first two terms of the expansion, thus obtaining the equation in the form
$L\frac{\partial I^{c}_{n}}{\partial t}=\frac{\partial^{2}V}{\partial Z^{2}}+\frac{1}{12}\frac{\partial^{4}V}{\partial Z^{4}}.$ (3)
(And, of course, in the framework of the approximation the derivative in (1a) should be understood as partial derivative.)
From (1a) and (3) we can obtain closed wave equation for the voltage $V$
$L\frac{\partial^{2}Q(V)}{\partial t^{2}}=\frac{\partial^{2}V}{\partial Z^{2}}+\frac{1}{12}\frac{\partial^{4}V}{\partial Z^{4}}.$ (4)
However, further on we will not use Eq. (4); it will be more convenient to use Eqs. (1a) and (3).
In this paper we are interested in the travelling wave solutions, for which the dependence of all the variables upon $Z$ and $t$ is of the form
@@
where $U$ is the speed of the travelling wave (the distance is measured in the units of the transmission line spacial period $\Lambda$). This gives us the opportunity to perform integration in (3) to get
$ULI^{c}=-\frac{\partial V}{\partial Z}-\frac{1}{12}\frac{\partial^{3}V}{\partial Z^{3}}.$ (6)
@@
$U^{2}L\frac{dQ}{dZ}=\frac{dV}{dZ}+\frac{1}{12}\frac{d^{3}V}{dZ^{3}}.$ (7)
Presenting $dQ/dz$ as $dQ/dz=(dQ/dV)(dV/dZ)$, introducing new dependent variable
@@
$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}=U^{2}L\frac{dQ(V)}{dV}-1.$ (10)
@@
$U^{2}L\left(\frac{dQ}{dV}\right)_{V=V_{1,2}}-1>0.$ (12)
The inequality (12) has a simple physical meaning. Consider the ”sound” waves that is small amplitude, smooth waves propagating on the homogeneous backgrounds $V=V_{1,2}$. These waves are described by Eq. (10), with the l.h.s. being put equal to zero. Hence the speed of the sound is given by the equation
$\frac{1}{u^{2}L}=\frac{dQ}{dV}.$ (13)
@@
The conditions (15) mean that the particle is at rest at the origin at ”time” equal to $V_{1}$ and then again at ”time” equal to $V_{2}$.
@@
$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}=U^{2}L\left(\frac{dQ}{dV}\right)_{V=V_{1,2}}-1=\text{const}.$ (16)
Taking into account the boundary conditions (15) we obtain
$E\sim\left(V-V_{1,2}\right),\hskip 28.45274pt\text{ in the vicinity of }V_{1,2}.$ (17)
@@
$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}=U^{2}L(\alpha+2\beta V+3\gamma V^{2})-1.$ (19)
For brevity we’ll call Eq. (18) and (19) the cubic case. Not that in this case
$\frac{1}{LC_{0}u^{2}(V)}=\alpha+2\beta V+3\gamma V^{2}.$ (20)
@@ II.2 The kinks
$V(Z)=\frac{V_{2}+V_{1}}{2}+\frac{V_{2}-V_{1}}{2}\tanh\left[\frac{V_{2}-V_{1}}{2}\psi Z\right].$ (27)
After finding $V(Z)$ we can from (2) obtain the current
$I(Z)=\frac{\psi\left(V_{2}-V_{1}\right)}{2UL}\tanh\left[\frac{V_{2}-V_{1}}{2}\psi Z\right].$ (28)
@@ II.3 The solitons
With three boundary conditions on $E(V)$ supplied by (15) and (29) for the soliton problem of the three parameters $V_{1},V_{3},U^{2}$ only two (say $V_{1}$ and $V_{3}$) can be considered as free parameters. Multiplying both sides of (7) by $V_{3}-V$, integrating the equation with respect to $V$ from $V_{1}$ to $V_{3}$ and taking into account the boundary conditions we obtain the speed of the soliton
@@
From (II.3) we see that the soliton exists both for positive and negative $\beta$. For positive $\beta$ we have dark soliton ($V_{3}<V_{1}$), and for negative $\beta$ – bright soliton ($V_{3}>V_{1}$). From (36c) we see that for any sign of $\beta$ the soliton speed is larger than the sound speed at the background, given by equation (20) with $\gamma=0$, in line with our previous general statement that the soliton is supersonic.
@@
Similar to what was done in the previous subsection, we can use (2) to find the current and obtain
$I(Z)=\frac{\phi\left(V_{3}-V_{1}\right)}{2UL}\operatorname{sech}^{2}\left(\frac{V_{3}-V_{1}}{2}\phi Z\right).$ (39)
@@ II.4 The quasi-continuum approximation
We would like to discuss an opportunity (and motivation) to go beyond the limits of the quasi-continuum approximation used in the paper. To do it, one may keep the third term in the Taylor expansion of the discrete second order derivative in the r.h.s. of (2) kogan2 ; kogan3
@@
$U^{2}L\frac{dQ}{dZ}=\frac{dV}{dZ}+\frac{1}{12}\frac{d^{3}V}{dZ^{3}}+\frac{1}{360}\frac{d^{5}V}{dZ^{5}}.$ (41)
@@ III The dissipative kinks and shocks
In this case (1c) is modified to
@@
hence Eq. (3) takes the form
$L\frac{\partial I^{c}}{\partial t}=\frac{\partial^{2}V}{\partial Z^{2}}+\frac{1}{12}\frac{\partial^{4}V}{\partial Z^{4}}-\frac{L}{R}\frac{\partial^{3}V}{\partial t\partial Z^{2}}.$ (43)
@@
From (1a) and (43) we can obtain Eq. (4) modified due to the presence of ohmic resistors
$L\frac{\partial^{2}Q(V)}{\partial t^{2}}=\frac{\partial^{2}V}{\partial Z^{2}}+\frac{L}{R}\frac{\partial^{3}V}{\partial t\partial^{2}Z}+\frac{1}{12}\frac{\partial^{4}V}{\partial Z^{4}}.$ (44)
For the travelling wave, combining (1a) and (43), we obtain after integration closed equation for $V(Z)$
$\frac{1}{12}\frac{d^{2}V}{dZ^{2}}+\frac{UL}{R}\frac{dV}{dZ}=U^{2}LQ-V+F,$ (45)
where $F$ is the constant of integration.
Equation (45) is Newton equation, describing motion with friction of the fictitious particle ($V$ being the coordinate of the particle and $Z$ the ”time”) in the force field given by the r.h.s. of the equation. We again impose on $V(Z)$ the boundary conditions (11). Hence the motion starts at $Z=-\infty$ at $V=V_{1}$ and ends at $Z=+\infty$ at $V=V_{2}$ (obviously that both $V_{1}$ and $V_{2}$ are the equilibrium positions). Equating the r.h.s. of (45) to zero at $V=V_{1}$ and $V=V_{2}$ we recover Eq. (21) for the speed of the wave.
It will be convenient to present Eq. (45) as
$\frac{1}{12}\frac{d^{2}V}{dZ^{2}}+\frac{UL}{R}\frac{dV}{dZ}=-\frac{d\Pi(V)}{dV},$ (46)
where the ”potential energy” $\Pi(V)$ is defined by the equation (which took into account the boundary conditions)
$\frac{d\Pi(V)}{dV}=-\frac{V_{2}-V_{1}}{Q_{2}-Q_{1}}Q+V+\frac{V_{2}Q_{1}-V_{1}Q_{2}}{Q_{2}-Q_{1}}.$ (47)
The constraint for the travelling wave asymptotic boundary conditions is the inequality
$\Pi(V_{1})>\Pi(V_{2}).$ (48)
Inequality (48) can be explicitly presented as
[…]
When both $\beta>0$ and $\gamma>0$, the condition (49) means just $|V_{1}|>|V_{2}|$.
The point $V=V_{1}$ should be the position of unstable equilibrium. The point $V=V_{2}$ can be the point of either stable, or unstable equilibrium. We met with similar dichotomy while studying travelling waves with finite support in Josephson transmission line kogan4 ). The stable equilibrium waves we called shocks, and the unstable equilibrium - kinks. There are several important distinctions between these two types of waves.
Unstable equilibrium at both at the point $V=V_{1}$ and $V=V_{2}$ means that differentiating the r.h.s. of (45) we recover equation (12). The kink, similar to what was in the lossless case, is supersonic both with respect to the region before the kink and the region behind the kink. Since the particle stops at the unstable equilibrium point, for the kink to exist, the fine tuning is necessary - the parameters $V_{1}$, $V_{2}$ and $R$ should satisfy definite relation.
Stable equilibrium at the point $V=V_{2}$ means
$U^{2}L\left(\frac{dQ}{dV}\right)_{V=V_{2}}-1<0.$ (50)
Inequalities (50) and (12) (the latter only for $V=V_{1}$) reflect the well-known in the nonlinear waves theory fact: the shock speed is larger than the sound speed in the region before the shock but smaller than the sound speed in the region behind the shock whitham .
To clarify the distinction between the shocks and the kinks let us bring together the formulas for the travelling wave speed and the sound speed for the cubic case
$\frac{1}{LC_{0}U^{2}}=\alpha+\beta\left(V_{1}+V_{2}\right)+\gamma\left(V_{1}^{2}+V_{1}V_{2}+V_{2}^{2}\right),$ (51a)
$\frac{1}{LC_{0}u^{2}(V_{1,2})}=\alpha+2\beta V_{1,2}+3\gamma V_{1,2}^{2}.$ (51b)
We realize that for $\beta>0$ and $\gamma>0$, for $V_{1}$ and $V_{2}$ having the same sign the travelling wave is always the shock. For $V_{1}$ and $V_{2}$ having opposite signs, we can have both the kink and the shock, depending upon the parameters.
To study quantitatively dissipative kinks and shocks we will use the approach of the previous Section. After repeating exactly what was done there we obtain modified Eq. (10)
$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}+\frac{UL}{R}\frac{dE}{dV}=U^{2}L\frac{dQ(V)}{dV}-1.$ (52)
The boundary conditions (11) lead to the boundary conditions (15) for $E(V)$, like it was in the lossless case. Equation (52) in the vicinity of $V_{1,2}$ has the form
$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}+\frac{UL}{R}\frac{dE}{dV}=\text{const},$ (53)
@@
In the cubic case Eq. (52) becomes
$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}+\frac{UL}{R}\frac{dE}{dV}=U^{2}L(\alpha+2\beta V+3\gamma V^{2})-1.$ (54)
We regain Eq. (23), and there exists kogan4 a simple polynomial solution of (54) in the form (24), which we repeat here
$E(V)=\psi(V-V_{1})(V-V_{2}),$ (55)
where $\psi$ is a constant (positive for positive $V_{1}$). Substituting (55) into (54) and equating the coefficients before the same powers of $V$ we obtain after simple algebra the generalization of (II.2)
@@
$V_{1}+V_{2}$ $=-\frac{2\beta}{3\gamma}+\frac{4UL}{\psi R}.$ (56b)
@@
Equation (55) shows that in the lossy line the kinks are of the same shape as in the lossless line (given by (27)). And the the shocks are of the same shape also. However, the value of $V_{2}$ (for a given $V_{1}$) depends upon $R$ (Eq. (56b) instead of (25b)) and so does the kink (the shock) speed, in spite of the validity in all the cases of Eq. (23).
The boundary in the phase space between the kinks and the shocks can be easily obtained from (III) in the case when $\beta=0$. Equating the speeds we obtain that this boundary corresponds to
$V_{1}+2V_{2}=0$ (57)
(very much similar to what we had for the Josephson transmission line kogan4 ).
We need to emphasise the different meaning of the condition (56b) for the kinks and the shocks. For the kinks, (56b) can be understood as the constraint on $V_{2}$ (for a given value of $V_{1}$ and ohmic resistance $R$). And it happens that the kink satisfying such constraint is described by the elementary function. For the shock, (56b) is just the condition for it being described by an elementary function. But, of course, the shock with any other value of $V_{2}$ (satisfying (48) also exists.
Let us assume that $V_{1}$ and $V_{2}$ are so close to each other, that we can keep in the r.h.s. of (18) only the first two terms, thus simplifying (54) to
$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}+\frac{UL}{R}\frac{dE}{dV}=U^{2}LC_{0}(\alpha+2\beta V)-1.$ (58)
The solution of (58) satisfying the boundary conditions (15) and (29) is
$E=\chi\left(V-V_{2}\right)\left[\left(V-V_{2}\right)^{1/2}-\left(\Delta V\right)^{1/2}\right],$ (59)
where $\Delta V\equiv V_{1}-V_{2}$ (we assumed for the sake of definiteness that $V_{1},V_{2}>0$). Substituting (59) into (58) we obtain
$\chi^{2}$ $=8U^{2}LC_{0}\beta$ (60a)
$\frac{UL}{R}$ $=\frac{5\chi}{24}\left(\Delta V\right)^{1/2}.$ (60b)
The function $E(V)$ being found, we can find the function $V(Z)$ by solving equation
$\frac{dV}{dX}=\chi\left(V-V_{2}\right)\left[\left(V-V_{2}\right)^{1/2}-\left(\Delta V\right)^{1/2}\right].$ (61)
The solution is
$V(Z)=V_{2}+\frac{\Delta V}{\left[\exp\left(\chi Z/2\right)+1\right]^{2}}.$ (62)
According to what was said after Eq. (III), Eq. (62) describes the shock. And we have to repeat here what was said after Eq. (III). Equation (60b) is just the condition for the shock to be expressed by an elementary function. The shock exists for any other value of $V_{2}$.